The revenue for a sandwich shop is directly proportional to its advertising budget. When the owner spent $2000 a month on advertasing, the revenue was $120,000. If the revenue is now 180,000, how much si the owner spending on advertising?
Answer by checkley75(3666)

2,000/120,000=x/180,000 now cross multiply:
120,000x=2,000*180,000
120,000x=360,000,000
x=360,000,000/120,000
x=$3,000 amount of advertising spent to increase the revenue to $180,000